OpenStudy (anonymous):

In pensacola in June, High tide was at noon. the water level at high tide was 12 feet and 2 feet at low tide.Assuming the next high tide is exactly 12 hours later and that the height of the water can be modeled by a cosine curve , find an equation for water level in june for pensacola as a function of time (t)

OpenStudy (anonymous):

@Preetha @praxer @cwrw238

OpenStudy (anonymous):

the possible choices are : a.) f(t)=12cos pi/2t+5 b.)f(t)=5 cos pi/2t+12 c.) f(t)=5 cos pi/6t+7 d.)f(t)=7cos pi/6t+12

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

the correct answer was c